OpenStudy (anonymous):

How fast is the dog moving when its kinetic energy has been reduced by half?

OpenStudy (anonymous):

we know that k=MV^2=here mass is constant then v^2=2k=0.5*0.5/2=0.125(ans)

OpenStudy (anonymous):

The kinetic energy is proportional to the velocity squared. When this is halved, the velocity is reduced to sqrt(0.5)=0.707 of its original value, 71% roughly.